CONTINUOUS-WAVE CAVITY RINGDOWN STUDY OF THE FIRST POSITIVE BAND SYSTEM OF N_{2}^{∗}

\underline{BRETT A. McGUIRE}, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; SUSANNA L. WIDICUS WEAVER^{a}, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801; MICHAEL WICZER, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; BOGDAN NEGRU, Department of Chemistry, University of California, Berkeley, CA 94720; JOSHUA P. DiGANGI, Department of Chemistry, University of Wisconsin, Madison, Wisconsin 53706; BRIAN A. TOM, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

We previously reported the assignment of the N_{2}^{+} A^{2}Π_{u} - X^{2}Σ_{g}^{+} 2-1 vibronic band from continuous-wave cavity ringdown spectroscopy (cw-CRDS) of an N_{2} plasma.^{b} More than 400 lines from the N_{2}^{∗} B^{3}Π_{g}-A^{3}Σ_{u}^{+} system were also observed during this experiment. While assignments for 37 vibronic bands from the first positive band system of N_{2}^{∗} have been reported in the literature, no assignments have been made within the frequency range of our laser. In this experiment, N_{2}^{∗} was formed in a positive column discharge of a mixture of N_{2} and argon, and spectra were collected between 937 and 984 nm. We have now identified lines from the 3-3 band of N_{2}^{∗}, and this spectral range is also expected to contain lines from the 2-2, 4-4, 7-8, and 8-9 bands. We will report on the progress of N_{2}^{∗} line assignment and the molecular constants determined from this work.

aCurrent address: Department of Chemistry, Emory University, Atlanta, GA 30322 bWidicus Weaver et al., J. Mol. Spec, 2008, in press.